A metal rod 80cm long lengthens by 0.090cm when its temperature rises by 93.6$^o$C. What is the linear expansivity of the metal

(a)0.0012/k
(b)0.000012/k
(c)12/k
(d)1.2/k
1
Expert's answer
2018-11-16T10:39:09-0500

Δl=l_0 αΔT

The linear expansivity

α=Δl/(l_0 ΔT)=0.090/(80×93.6)=1.2×〖10〗^(-5)=0.000012K^(-1)

(b) 0.000012K^(-1)